Find the mean of the probability distribution of a random variable X which if 𝑃(𝑋) =

đ‘„+1

20

for X= 1, 2, 3, 4, and 5.


1
Expert's answer
2022-03-24T09:45:41-0400

"P(X=1)={\frac 2 {20}}"

"P(X=2)={\frac 3 {20}}"

"P(X=3)={\frac 4 {20}}"

"P(X=4)={\frac 5 {20}}"

"P(X=5)={\frac 6 {20}}"

"M(W)={\frac 2 {20}}*1+{\frac 3 {20}}*2+{\frac 4 {20}}*3+{\frac 5 {20}}*4+{\frac 6 {20}}*5={\frac {70}{20}}=3.5"